### v2

# Can MLLMs Guide Me Home? A Benchmark Study on Fine-Grained Visual Reasoning from Transit Maps

Abstract
Multimodal large language models (MLLMs) have recently achieved significant progress in visual tasks, including semantic scene understanding and text-image alignment, with reasoning variants enhancing performance on complex tasks involving mathematics and logic. However, their capacity for reasoning tasks involving fine-grained visual understanding remains insufficiently evaluated. To address this gap, we introduce ReasonMap, a benchmark designed to assess the fine-grained visual understanding and spatial reasoning abilities of MLLMs. ReasonMap encompasses high-resolution transit maps from $30$ cities across $13$ countries and includes $1,008$ question-answer pairs spanning two question types and three templates. Furthermore, we design a two-level evaluation pipeline that properly assesses answer correctness and quality. Comprehensive evaluations of $15$ popular MLLMs, including both base and reasoning variants, reveal a counterintuitive pattern: among open-source models, base models outperform reasoning ones, while the opposite trend is observed in closed-source models. Additionally, performance generally degrades when visual inputs are masked, indicating that while MLLMs can leverage prior knowledge to answer some questions, fine-grained visual reasoning tasks still require genuine visual perception for strong performance. Our benchmark study offers new insights into visual reasoning and contributes to investigating the gap between open-source and closed-source models.
1 Introduction
Figure 1: Overview of ReasonMap. ReasonMap is a benchmark dataset designed to evaluate fine-grained visual reasoning abilities of MLLMs, encompassing $1,008$ question–answer pairs constructed over high-resolution transit maps from $30$ cities, spanning two question types and three templates.
Multimodal large language models (MLLMs) [1, 2, 3, 4, 5] have recently achieved notable advancements across a range of vision-language tasks, including visual grounding [6, 7], reasoning segmentation [8, 9, 10, 11], and text-image alignment [12, 13]. Building upon these developments, reasoning MLLMs [14, 15, 16, 17, 18, 19, 20] have further improved performance on complex visual reasoning tasks such as visual math problems [21, 22], visual question answering (VQA) [23, 24, 22], and spatial reasoning [23, 25, 26]. These capabilities are critical for a wide range of real-world applications, including embodied AI, autonomous agents, and decision-making systems such as autonomous driving [27, 28, 29]. As multimodal tasks grow in complexity and practical relevance, the need for rigorous benchmarks to assess fine-grained visual reasoning becomes increasingly essential.
To address the growing demand for robust evaluation of multimodal reasoning, several benchmarks have been proposed. Datasets such as MathVQA [22] and MMMU [24] incorporate multimodal questions but often permit models to succeed via shallow heuristics, without requiring genuine visual grounding. MathVerse [30] mitigates this limitation by introducing diverse problem variants that encourage reliance on visual input. VisuLogic [31] further enforces visual reasoning by explicitly eliminating language-only shortcuts. Other efforts, such as VisualPuzzles [32], VGRP-Bench [33], and R-Bench [34], target logical and structural reasoning, while CityBench [35] and DriveBench [36] focus on domain-specific applications like urban tasks and autonomous driving. VBench [37] emphasizes detailed visual understanding. MapBench [38] addresses spatial reasoning by introducing structured scene graphs for map navigation. Despite these advances, systematic evaluation of fine-grained visual reasoning remains limited, especially for structured and information-rich diagrams like high-resolution transit maps, leaving a critical gap in existing benchmarks.
In this paper, we introduce ReasonMap (Figure 1), a benchmark designed to evaluate the fine-grained visual understanding and spatial reasoning capabilities of MLLMs using high-resolution transit maps. As structured and information-dense visual artifacts, transit maps inherently require precise spatial interpretation, making them well-suited for assessing detailed visual reasoning. ReasonMap comprises $1,008$ human-verified question-answer pairs spanning $30$ cities across $13$ countries. Each instance includes a map, two stops, two questions (short and long), multiple reference routes, and two difficulty labels (map and question difficulty). The questions cover two types and three prompting templates—one for short and two for long questions—capturing both coarse and fine-grained spatial reasoning. To ensure data quality, we perform manual route verification, promote question diversity, and balance difficulty distribution. For evaluation, we propose a two-level framework that independently measures answer correctness (via accuracy) and quality (via a proposed map score), reflecting both feasibility and efficiency in route planning.
We conduct comprehensive experiments on $15$ widely-used MLLMs, encompassing base and reasoning models. Our results reveal a counterintuitive finding: among open-source models, base variants outperform their reasoning counterparts, whereas the opposite holds for closed-source models. Moreover, when only textual inputs are provided, models can still answer some questions based on inner knowledge, but in most cases, their performance noticeably drops. This highlights a critical limitation in the current model behavior. While some models can leverage prior knowledge and textual cues to solve certain tasks, the tasks (e.g., fine-grained visual reasoning tasks) requiring genuine visual understanding still necessitate effective integration of multimodal information for robust reasoning.
Our main contributions are summarized as follows:
  • We develop an extensible, semi-automated pipeline for dataset construction, facilitating scalable expansion to additional maps and cities. Using this pipeline, ReasonMap is constructed to evaluate fine-grained visual reasoning capabilities in MLLMs.
  • We propose a structured two-level evaluation framework that separately quantifies answer correctness and quality using accuracy and the proposed map score, respectively, enabling fine-grained answer assessment.
  • A comprehensive benchmarking study is conducted across $15$ MLLMs, providing insights into model performance, robustness, and the interplay between visual and textual cues, thereby informing future research on multimodal reasoning.
2 Related Work
Reasoning in LLMs & MLLMs. Recent advances in large language models (LLMs) have demonstrated significant improvements in reasoning capabilities through reinforcement fine-tuning paradigms [14, 15, 39, 40], which leverage GRPO [41] to unlock the reasoning potential of LLMs. This paradigm has also been extended to the multimodal domain, with increasing interest in applying reinforcement learning (RL) to visual reasoning [42, 43, 44, 45, 46]. Both open-source and closed-source communities have introduced advanced reasoning MLLMs built upon earlier systems [3, 2, 47, 48]. Notable open-source models include Kimi-VL [16], Skywork-R1V [17, 18], and Qwen-QVQ [20], while Doubao-1.5-Pro [19], OpenAI o3 [48], OpenAI 4o [49], and Gemini [50] represent leading closed-source efforts. Despite recent progress, systematic evaluation of fine-grained visual reasoning in MLLMs remains limited, as existing benchmarks primarily target coarse-grained tasks and fail to capture model performance on complex real-world visual content.
Table 1: Comparison between ReasonMap and existing multimodal reasoning datasets. For entries in the dataset size column with notation like (×$\times n$), each base problem has multiple versions to enforce visual grounding. Specifically, VGRP-Bench is constructed by sampling over 20 core puzzles.
Name Year Dataset Size Avg. Resolution Training Set Step Evaluation Multilingual (Count)
MMMU [24] 2024 $11.5$k $684\times 246$ ($2$)
MathVerse [30] 2024 $2,612$ (×$\times 6$) $577\times 487$
VisuLogic [31] 2025 $1,003$ $601\times 331$
VisualPuzzles [32] 2025 $1,168$ $767\times 464$
VGRP-Bench [33] 2025 $20$ (×$\times 5$) $790\times 790$
R-Bench [34] 2025 $665$ $629\times 348$ ($2$)
VBench [37] 2023 $191$ $2,246\times 1,582$
ReasonMap 2025 $1,008$ (×$\times 2$) $5,839\times 5,449$ ($4$)
Multimodal Reasoning Datasets. As multimodal reasoning has rapidly progressed, various benchmarks have emerged to evaluate MLLMs across different reasoning dimensions (see summary in Table 1). Datasets such as VBench [37], VisualPuzzles [32], VisuLogic [31], and VGRP-Bench [33] primarily examine abstract visual reasoning through synthetic tasks involving logic, structure, and pattern recognition. In parallel, CityBench [35] and DriveBench [36] shift focus to real-world spatial reasoning, assessing model performance on complex urban or autonomous driving scenarios. For mathematical reasoning, benchmarks like MathVQA [22], MMMU [24], and MathVerse [30] integrate multimodal inputs, with MathVerse notably introducing varied problem formats to strengthen visual dependence. Additionally, MapBench [38] employs structured scene graphs combined with CoT prompting to support navigation tasks based on manually curated and verified questions. Its image resolution, while relatively low, reflects a common characteristic shared by current datasets. Unlike these works, we first introduce a benchmark for evaluating fine-grained visual reasoning capacities with high-resolution transit maps.
Map-based Spatial Reasoning. Among the many directions of multimodal reasoning, map-based spatial reasoning has emerged as a crucial area, with broad applications in navigation, urban planning, and autonomous systems [51, 52, 53, 54]. Recent efforts have focused on enabling models to interpret and reason over various types of map data. CityBench [35] provides a dataset for evaluating urban scene understanding, while MapLM [55] introduces a benchmark for map and traffic scene understanding. PlanAgent [56] and PReP [57] explore embodied planning in environments that require interpreting map information. MapEval [26] proposes a structured evaluation suite for map-related reasoning, and GeoNav [53] investigates geospatial navigation using LLMs. Most existing methods [26, 35, 56] depend on external tools (e.g., map services or APIs) to complete spatial tasks, which often bypasses the need for genuine visual reasoning. However, spatial reasoning based solely on visual understanding remains essential. Our work targets to fill this gap by evaluating such capabilities without tool assistance.
3 ReasonMap Construction
Figure 2: The building pipeline of ReasonMap consists of three main stages: (1) data collection and preprocessing, (2) question–answer pair construction, and (3) quality control. Steps (2-4) in the figure correspond to the question–answer pair construction stage.
In this section, we first present the complete dataset building pipeline as shown in Figure 2, which consists of three main stages: (1) data collection and preprocessing; (2) question–answer pair construction; and (3) quality control. We then report comprehensive statistics of the dataset.
3.1 ReasonMap Building Pipeline
3.1.1 Data Collection and Preprocessing
We collect and manually select $30$ high-resolution transit maps covering $30$ cities across $13$ countries from publicly available online sources, in compliance with relevant licenses and regulations, ensuring diversity and a balanced range of map difficulty. We then leveraged MLLMs to extract the names of transit lines and their corresponding stops, followed by manual correction to ensure correctness. Special cases such as transfer stops and branch-starting stops were annotated in a standardized format appended to the respective stop entries. Finally, for subsequent usage, all route and stop information was saved in a unified JSON format, referred to as the Map Metadata.
3.1.2 Question-Answer Pair Construction
The construction of question–answer pairs involves three key steps: (1) Question Generation, where we formulate questions based on predefined templates; (2) Reference Route Collection, where we obtain corresponding reference routes using Gaode Map [^1] https://console.amap.com/dev/index and Google Map [^2] https://developers.google.com/maps/apis-by-platform; and (3) Label Annotation, where we assign difficulty levels for both the maps and the questions.
Question Generation. We randomly select two stops (refer to stop1 and stop2) from the current high-resolution transit map. We then generate one short question and one long question based on predefined question templates and two stops (Figure 2). The short question has only one fixed template, while the long question is randomly assigned one of two available templates during generation. Additionally, the two long question templates differ in focus: one asks for the number of via stops, while the other requires identifying each via stop (see detailed templates in Appendix A.1).
Reference Route Collection. For each question, we query all valid transit routes between stop1 and stop2 using APIs from map services (e.g., Gaode Map for Chinese cities and Google Map for other cities). The retrieved routes are stored in a unified format containing relevant metadata (e.g., route name, departure stop, arrival stop, via stops, and number of via stops). We discard routes that cannot be visually traced on the map, ensuring consistency with the visual content.
Label Annotation. Two levels of difficulty labeling are included in this stage. For map difficulty, we manually assign each map to one of three difficulty levels (easy, medium, hard), ensuring a balanced split across $30$ maps, with $10$ maps per level. For question difficulty, we assign difficulty based on the number of transfers in the reference route: routes with no transfers are labeled as easy, those with one transfer as medium, and all others as hard. To ensure balance, we set a fixed difficulty distribution threshold of $20:15:5$ (easy:medium:hard) for each map, generating $40$ questions. Once the quota for a difficulty level is reached on a given map, no additional questions of that level are retained.
3.1.3 Quality Control
To ensure the reliability and balance of the dataset, we perform quality control from three perspectives: correctness, diversity, and difficulty balance. Incorrect question–answer pairs are either manually corrected or discarded. We then involve both automatic checks and manual adjustments to ensure consistency and coverage across all difficulty levels. One reserved example is shown in Figure 1.
3.2 Dataset Statistics
The ReasonMap consists of $30$ high-resolution transit map images (see map sources in Appendix A.2) with an average resolution of $5,839\times 5,449$ pixels. In total, it contains $1,008$ question–answer pairs, including stop names in four languages (e.g., English, Hungarian, Chinese, and Italian). The distribution of question difficulty is as follows: $57.7\%$ are labeled as easy, $34.4\%$ as medium, and $7.8\%$ as hard. Additionally, a subset of $312$ samples is manually selected as the test set for the benchmark experiments described in Section 5, while the remaining samples serve as a training set for future use. To ensure diversity and difficulty balance, the test set includes $11$ cities with a $4:3:4$ map difficulty ratio and a question difficulty distribution ($181$ easy, $108$ medium, $23$ hard) that maintains consistency with the full dataset.
4 Evaluation Framework
This section systematically introduces a two-level evaluation framework for assessing model performance on the ReasonMap. This framework separately evaluates the correctness and quality of answers produced by models. Specifically, we quantify correctness using accuracy and design map score to measure the quality of answers, considering multiple factors (e.g., route efficiency and alignment with the reference routes from map services).
Preparation for Evaluation. We first parse the model-generated answers according to the required format. Answers that do not comply with the specified format or cannot be parsed due to model hallucination [58] are marked as invalid. Invalid responses are excluded from subsequent evaluations, with accuracy and map score set to zero. For the correctness evaluation, we utilize the Map Metadata mentioned in Section 3.1.1 as ground truth. For the quality evaluation, we adopt the collected reference routes as presented in Section 3.1.2 as the ground truth.
4.1 Correctness Evaluation
We evaluate the correctness of the answer using Algorithm 1 in Appendix B. Specifically, the evaluation checks the correctness of the overall departure and arrival stops (stop1 and stop2), verifies if the route name of each segment exists in the Map Metadata, ensures the departure and arrival stops are valid for each segment, and confirms that transfer stops between consecutive segments are consistent. An answer is considered correct only if all the above checks are satisfied. Additionally, we apply the same correctness evaluation algorithm to the answers of short and long questions.
4.2 Quality Evaluation
To evaluate the quality of the answers, we introduce a unified scoring metric, referred to as the map score, which is applied to both short and long questions using the evaluation procedure (see Algorithm 2 in Appendix B). The overall evaluation framework for route quality follows a structure similar to that used in Section 4.1. The following evaluation procedure assumes a single reference route for simplicity. In practice, if multiple reference routes are available, the answer is evaluated against each of them, and the highest score is taken as the final map score.
For short questions, the map score solely focuses on route-level and endpoint consistency, excluding all long-question-specific parts. We compute the score by comparing segment pairs in the answer and reference route. Specifically, correctly matching stop1 and stop2 contributes one point, matching the route name adds two points, and matching the departure and arrival stops within each route segment provides one point each. The final score is capped at $10$, and an additional bonus is awarded if the answer is judged correct based on the correctness evaluation procedure described in Section 4.1. This design ensures that a correct answer always receives a higher score than any incorrect one.
For long questions, the evaluation incorporates additional scoring components tailored to the two question templates introduced in Section 3.1.2. These components are designed to capture the increased reasoning depth required in long-form responses. As with short questions, a bonus score is also added for correct answers. The two additional scoring components are detailed below.
Via Stop Count Evaluation. For long questions that require models to predict the number of via stops for each segment, we introduce the num_via_stop_score. This score compares the via stop count of the answer and reference route by computing the absolute error and mapping it to a fixed score ($4$). A perfect match yields full points, while larger discrepancies receive proportionally lower scores. The score is then capped at $10$ for the full route.
Specific Via Stop Evaluation. For long questions that require explicit enumeration of intermediate stops, we compute via_stop_score using a combination of two factors: the number of correctly matched via stops, and the intersection-over-union (IoU) between via stop sets of the answer and reference route. The final score for this component is obtained by averaging the IoU score (scaled to $10$) and the exact match count (capped at $10$), and then clipped to a maximum of $10$.
5 Experiments
Table 2: Evaluations of various MLLMs on ReasonMap. $S.$ represents results for short questions, while $L.$ denotes results for long questions. The map score is capped at 20 for short questions, while for long questions, the maximum score is 40. Bold indicates the best results among open-source and closed-source models, respectively, while underline represents the second best.
Model Type Weighted Acc. ($S.$) #Tokens ($S.$) Weighted Acc. ($L.$) #Tokens ($L.$) Weighted Map Score ($S.$ / $L.$)
Open-source Models
Qwen2.5-VL-3B-Instruct [2] Base $8.68\%$ $42$ $7.99\%$ $151$ $2.75$ / $3.70$
Qwen2.5-VL-32B-Instruct [2] Base $16.49\%$ $36$ $15.71\%$ $112$ $3.88$ / $6.84$
Qwen2.5-VL-72B-Instruct [2] Base $\mathbf{26.65\%}$ $33$ $\mathbf{24.22\%}$ $104$ $\mathbf{5.09}$ / $\mathbf{8.80}$
InternVL3-38B [3] Base $14.84\%$ $43$ $13.45\%$ $68$ $3.48$ / $6.31$
InternVL3-78B [3] Base $\underline{25.35\%}$ $33$ $\underline{19.62\%}$ $62$ $\underline{4.80}$ / $\underline{7.50}$
Kimi-VL-A3B-Instruct [16] Base $12.76\%$ $41$ $12.33\%$ $41$ $3.30$ / $5.37$
Kimi-VL-A3B-Thinking [16] Reasoning $5.47\%$ $754$ $5.47\%$ $1,287$ $2.44$ / $3.17$
Skywork-R1V-38B [17, 18] Reasoning $6.86\%$ $645$ $3.21\%$ $842$ $2.11$ / $3.11$
QvQ-72B-Preview [20] Reasoning $9.03\%$ $1,279$ $4.25\%$ $1,619$ $1.59$ / $1.55$
Closed-source Models
Doubao-115 [19] Base $34.20\%$ $32$ $38.02\%$ $118$ $5.25$ / $11.96$
OpenAI 4o [49] Base $41.15\%$ $34$ $42.80\%$ $58$ $6.84$ / $13.57$
Doubao-415 [19] Reasoning $43.14\%$ $536$ $\underline{46.09\%}$ $1,796$ $7.33$ / $\underline{14.67}$
Doubao-428 [19] Reasoning $37.15\%$ $532$ $37.85\%$ $2,167$ $5.52$ / $11.73$
Gemini-2.5-Flash [50] Reasoning $\underline{46.09\%}$ $806$ $29.86\%$ $1,419$ $\underline{7.64}$ / $9.98$
OpenAI o3 [48] Reasoning $\mathbf{63.02\%}$ $1,236$ $\mathbf{59.11\%}$ $2,372$ $\mathbf{9.53}$ / $\mathbf{17.96}$
5.1 Experimental Setups
We conduct extensive benchmark experiments on ReasonMap using $15$ popular MLLMs under different inference settings, analyzing their performance and comparing results. Several interesting insights emerge from this comparison. The detailed experimental settings are described below.
Evaluated Models. We evaluate a diverse set of MLLMs categorized into two groups based on whether they are reasoning-oriented models with a long-thinking process. Reasoning models include: Skywork-R1V-38B [17, 18], QvQ-72B-Preview [20], Kimi-VL-A3B-Thinking/Instruct [16], OpenAI o3 [48], Gemini-2.5-Flash [50], Doubao-1-5-thinking-vision-pro-250428 (Doubao-428), and Doubao-1.5-Thinking-Pro-M-250415 (Doubao-415) [19]. Base models include: Qwen2.5-VL series (3B, 32B, 72B) [2], InternVL3 series (38B, 78B) [3], OpenAI 4o [49], and Doubao-1.5-Vision-Pro-32k-250115 (Doubao-115) [19]. Additionally, the Doubao 1.5 Pro series has an activated parameter size of $20$B.
Inference Settings. For open-source models, we set the max output token limit to $2,048$, while keeping other parameters consistent with the official HuggingFace configurations. All open-source models are deployed using PyTorch and the HuggingFace Transformers library on $8$ NVIDIA A100 GPUs. For closed-source models, we access their official APIs for evaluation and follow the default settings provided by each model’s official documentation. We further discuss the diverse image processing strategies when handling high-resolution visual inputs in Appendix D.
Difficulty-Aware Weighting. To better reflect the varying complexity of different samples, we adopt a difficulty-aware weighting strategy based on the combination of question difficulty and map difficulty. Specifically, each difficulty pair is assigned a predefined weight, with harder combinations receiving higher values. The complete weight matrix is provided in Appendix B.2. Both accuracy and map score are evaluated using this weighting scheme, ensuring that models are more strongly rewarded for correctly solving more challenging examples.
5.2 Experimental Results
5.2.1 Performance of MLLMs with Full Input
Refer to caption Refer to caption
(a) Accuracy for short questions (b) Accuracy for long questions
Figure 3: Accuracy across difficulty combinations for four representative MLLMs (Qwen2.5-VL-72B-I, InternVL3-78B, OpenAI o3, and Doubao-415). Each difficulty combination is denoted by a pair (e.g., easy-hard), where the first term indicates question difficulty and the second term represents map difficulty. The pair (hard-middle) contains only one sample, leading to an accuracy of 100%. We summarize the number of evaluation samples in each difficulty bucket: $55$ samples for easy-easy, $46$ for easy-middle, $28$ for middle-easy, $7$ for hard-easy, $23$ for middle-middle, $80$ for easy-hard, $1$ for hard-middle, $57$ for middle-hard, and $15$ for hard-hard.
The principal results are summarized in Table 2. Notably, we observe a counterintuitive phenomenon: among open-source models, reasoning models consistently underperform their base counterparts, whereas the opposite holds in the closed-source setting [^3] Although the comparison across closed-source models may not be fair due to lack of transparency in details, the reasoning variants exhibit stronger performance in this category.. Prior work suggests that RL may improve sample efficiency without introducing fundamentally new reasoning abilities [59, 60, 61], while RL-trained models tend to bias their output distributions toward high-reward trajectories, which helps produce more correct responses but may simultaneously constrain the model’s exploration capacity and reduce its ability to leverage broader foundational knowledge. In addition, recent studies indicate that multimodal models may sometimes rely on inner knowledge priors instead of truly attending to visual inputs [62, 63, 64, 30]. This tendency is further supported by the results in Section 5.2.2, where open-source models still maintain part of their performance even without visual input, indicating limited visual grounding. In contrast, closed-source reasoning models outperform their base variants. One possible explanation lies in the broader knowledge coverage and better visual integration observed in these models [19, 48, 50].
We further analyze the effect of model size by examining performance within the same architecture series. Qwen2.5-VL and InternVL series show a consistent trend: larger models achieve better accuracy with fewer tokens, suggesting that the scaling law [65] continues to hold even in fine-grained visual reasoning tasks. Figure 3 presents accuracy distributions across different combinations of question and map difficulty. As expected, performance degrades as task complexity increases. Additionally, Figure 4 illustrates accuracy variation across cities. We observe a negative correlation between map difficulty and accuracy. Moreover, model performance varies notably even among cities with comparable map difficulty levels. This disparity can be partially attributed to factors such as city prominence and the language used for stop names, both of which are closely tied to the model’s pretrained knowledge. For instance, OpenAI o3 performs significantly better on complex cities like Singapore compared to Hangzhou, likely due to Singapore’s higher international visibility and the use of English stop names, whereas Hangzhou is less prominent and its stop names are Chinese.
Refer to caption Refer to caption
(a) Accuracy for short questions (b) Accuracy for long questions
Figure 4: Accuracy across different cities for four representative MLLMs (Qwen2.5-VL-72B-I, InternVL3-78B, OpenAI o3, and Doubao-415). Each city is marked with the corresponding map difficulty and the country flag. Each city in the test set provides a specific number of samples per model: $32$ samples for Auckland, $34$ for Los Angeles, $7$ for Miami, $35$ for Lisboa, $18$ for Geneva, $40$ for Beijing, $39$ for Hangzhou, $17$ for Budapest, $39$ for Singapore, $40$ for Rome, and $11$ for Toronto.
Table 3: Evaluations of various MLLMs on ReasonMap without visual inputs. $S.$ represents results for short questions, while $L.$ denotes results for long questions. The map score is capped at 20 for short questions, while for long questions, the maximum score is 40. Bold indicates the best results among open-source and closed-source models, respectively, while underline represents the second best. Green highlights improved results compared to the full input setting (Table 2), while red indicates performance drops.
Model Type Weighted Acc. ($S.$) #Tokens ($S.$) Weighted Acc. ($L.$) #Tokens ($L.$) Weighted Map Score ($S.$ / $L.$)
Open-source Models
[…]
Closed-source Models
[…]
5.2.2 Performance of MLLMs without Visual Input
To further investigate the reliance of MLLMs on visual input, we selected representative open-source and closed-source models for additional experiments, where the visual input was masked. The results are reported in Table 3. We observe that while most models can leverage internal knowledge to answer certain questions, their performance generally declines to varying degrees when visual input is removed, with the decline being more pronounced among closed-source models. Model performance is positively correlated with the performance drop after masking visual inputs, indicating effective use of visual information. In contrast, models like Qwen2.5-VL-3B-I show minimal or even improved performance, suggesting a reliance on inner knowledge rather than genuine visual reasoning.
5.3 Error Analysis
Figure 5: Error case analysis of various MLLMs using ReasonMap. For reasoning models, the reasoning process is explicitly marked with <think> and </think> tags. We highlight error contents in the answers with red and categorize them accordingly.
Figure 5 presents representative failure cases from ReasonMap, revealing several recurring error types. A common issue is visual confusion, where the model misidentifies the transit line due to similar colors or adjacent layouts, for instance, mistaking Line $9$ for Line $16$ (OpenAI o3, left column; Doubao-428, right column). Another frequent problem is format errors, where responses deviate from the required structure, making them unprocessable despite containing correct route information (Doubao-115 and QvQ-72B-Preview, left column). We also observe instances of hallucination [58], where the model repeats the correct answer (Kimi-VL-A3B-Thinking, middle column) or generates information that is not present in the input, such as mentioning image generation, as seen in Skywork-R1V-38B (right column). Refusal cases are also present, where models explicitly decline to answer (Skywork-R1V-38B, middle and right column). Notably, these errors may occasionally co-occur within a single response (Skywork-R1V-38B, right column). These behaviors highlight the limitations in visual grounding and response robustness, especially when handling fine-grained visual details (see more case analyses in Appendix C).
6 Conclusion
In this work, we introduce ReasonMap, a benchmark designed to evaluate the fine-grained visual understanding and spatial reasoning capabilities of MLLMs using high-resolution transit maps. Through a semi-automated and scalable data building pipeline, we curate a diverse set of human-verified question-answer pairs across $30$ cities. Our two-level evaluation framework enables a nuanced assessment of both correctness and quality. Experimental results on $15$ popular MLLMs reveal key insights into model behavior, highlighting performance gaps between base and reasoning models, as well as the crucial role of visual input. These findings underscore the need for more rigorous evaluation and training approaches to advance visual reasoning in multimodal systems.
Appendix
We provide a comprehensive overview in the Appendix, covering key components of our dataset, methodology, and analysis. Specifically, we include the question templates used during dataset construction, sources of transit maps from $30$ cities, detailed descriptions of the evaluation algorithm, experimental setup, and in-depth analyses of both correct and erroneous cases.
In addition, we further discuss diverse image processing strategies employed by various MLLMs when handling high-resolution visual inputs. We also address the stated limitations of our work, potential broader impacts, ethical considerations related to the dataset, and licensing information for the MLLMs used in our experiments.
[…]
Appendix A Dataset Construction Details
A.1 Question Template Summary
We present one short question template and two long question templates as follows.
Short Question Template According to the subway map, how do I get from [Stop 1] to [Stop 2]? Provide only one optimal route, with only the line name and the departure and arrival stations. The format should be strictly followed:
Route Name: Line x
Departure Stop: xx Station
Arrival Stop: xx Station
Route Name: Line x
Departure Stop: xx Station
Arrival Stop: xx Station
Long Question Template 1 According to the subway map, how do I get from [Stop 1] to [Stop 2]? Provide only one optimal route, and include the number of via stops for each route section (excluding the departure and arrival stops). The format should be strictly followed:
Route Name: Line x
Departure Stop: xx Station
Arrival Stop: xx Station
Number of Via Stops: x
Route Name: Line x
Departure Stop: xx Station
Arrival Stop: xx Station
Number of Via Stops: x
Long Question Template 2 According to the subway map, how do I get from [Stop 1] to [Stop 2]? Provide only one optimal route, including all the via stops. The format should be strictly followed:
Route Name: Line x
Departure Stop: xx Station
Arrival Stop: xx Station
Via Stops: xx Station, xx Station, xx Station
Route Name: Line x
Departure Stop: xx Station
Arrival Stop: xx Station
Via Stops: xx Station
A.2 Map Source
We provide the sources of all maps included in ReasonMap for further reference (Table A1).
Table A1: Source links for city transit maps used in ReasonMap. We present a total of $30$ cities sourced from $13$ countries.
City Source City Source City Source
Budapest Link Oslo Link Rome Link
Lisboa Link Geneva Link Dubai Link
Auckland Link Sydney Link Singapore Link
Kuala Lumpur Link Los Angeles Link Miami Link
New York Link Toronto Link Washington Link
Guiyang Link Shanghai Link Huhehaote (Hohhot) Link
Nanchang Link Nanning Link Shenzhen Link
Hangzhou Link Dalian Link Kunming Link
Hefei Link Beijing Link Changzhou Link
Jinan Link Xi’an Link Changshang Link
Appendix B Evaluation Details
B.1 Correctness and Quality Evaluation
We present the detailed algorithms for evaluating answer correctness and quality in Section 4 (Algorithm 1 for correctness evaluation and Algorithm 2 for quality evaluation).
Initialize acc $\leftarrow 1$;
if departure stop of first segment $\neq$ stop1 or arrival stop of last segment $\neq$ stop2 then
       acc $\leftarrow 0$;
foreach segment in predicted route do
       if route name not in the Map Metadata then
             acc $\leftarrow 0$;
      if departure or arrival stop not in the stop list of the route then
             acc $\leftarrow 0$;
      if not the last segment then
             if arrival stop of current segment $\neq$ departure stop of next segment then
                   acc $\leftarrow 0$;
return acc
Algorithm 1 Correctness Evaluation
Initialize map_score $\leftarrow 0$;
if departure stop of first segment = stop1 and arrival stop of last segment = stop2 then
[…]
      /* Long-question-specific part */
       Initialize $\mathcal{V}_{\text{union}}$, $\mathcal{V}_{\text{intersection}}\leftarrow\emptyset$;
       Initialize via_stop_score, num_via_stop_score $\leftarrow 0$;
      foreach segment pair (answer route, reference route) do
             if answer route name = reference route name then
[…]
            if answer departure stop = reference departure stop then
[…]
            if answer arrival stop = reference arrival stop then
[…]
            /* Long-question-specific part */
             Calculate absolute difference ($error$) in the number of via stops;
[…]
            if answer route name = reference route name then
                   Update $\mathcal{V}_{\text{union}}$, $\mathcal{V}_{\text{intersection}}$ with answer and reference via stops respectively;
            via_stop_score $\leftarrow$ via_stop_score + number of correctly matched via stops;
      /* Long-question-specific part */
[…]
      via_stop_score $\leftarrow$ average( $|\mathcal{V}_{\text{intersection}}|/|\mathcal{V}_{\text{union}}|\times 10$, via_stop_score)
        map_score $\leftarrow$ map_score $+$ Option(via_stop_score or num_via_stop_score);
/* $10$ for short question; $20$ for long question */
[…]
if correctness evaluation (acc) = 1 then
[…]
return map_score;
Algorithm 2 Quality Evaluation
B.2 Details about Difficulty-Aware Weighting.
Each difficulty pair is assigned a predefined weight that reflects its relative challenge level. The full weight matrix is shown below, where the first element in each pair denotes the question difficulty and the second denotes the map difficulty:
("easy", "easy"): $1.0$ ("middle", "easy"): $1.5$ ("hard", "easy"): $2.0$
("easy", "middle"): $1.5$ ("middle", "middle"): $2.0$ ("hard", "middle"): $2.5$
("easy", "hard"): $2.0$ ("middle", "hard"): $2.5$ ("hard", "hard"): $3.0$
This weighting scheme rewards models more for correctly solving harder question–map combinations, reflecting the increased reasoning complexity they entail, while maintaining moderate differences between buckets to prevent excessive score variance and preserve evaluation stability.
Appendix C Case Analysis
We provide additional case analyses covering both correct and incorrect predictions, along with detailed comparisons of their respective reasoning processes.
We compare Doubao-415 and Doubao-428 (Figure A1), both of which reach the correct destination (from Augustins Station to Poterie Station) but via distinct reasoning paths. Doubao-415 correctly identifies early that both stations are on Line 18 and efficiently converges on the optimal, direct route without transfers. In contrast, Doubao-428 misclassifies Augustins as being on Line 12 and, assuming Poterie is on Line 18, proposes a transfer route via Plainpalais—functionally correct but suboptimal due to unnecessary complexity. Both models engage in extensive self-correction ($7270$ tokens for Doubao-428; $4474$ for Doubao-415), highlighting the significant downstream impact of early-stage misjudgments. Moreover, visual reasoning limitations persist: despite correctly recognizing Augustins on Line 12, Doubao-415 commits to a transfer path and fails to re-evaluate the possibility of a direct connection. This indicates room for improvement in both early visual grounding and global route optimality awareness.
We then analyze the observed pattern when comparing the full input and text-only variants in the case (in Figure A2). The model with full visual access accurately identifies both stations on the Yellow Line and outputs the optimal direct route with the correct number of via stops. In contrast, the text-only variant makes an early misclassification, placing both stations on the Blue Line (Azul) and constructing a plausible but entirely incorrect sequence of intermediate stops. Although the final answer format appears coherent, the underlying logic is flawed due to the initial error in line recognition. This further illustrates the importance of visual input in spatial reasoning tasks, where even minor misinterpretations can lead to fundamentally incorrect conclusions. Additionally, some models, such as the InternVL3 series, default to rejection when visual input is absent.
We further present several error cases (in Figure A3), where Doubao-415 still exhibits visual confusion. In contrast, Qwen2.5-VL-32B-I, when lacking visual input, behaves differently from the InternVL3 series: rather than rejecting the query outright, it attempts to reason over the available information without producing a final answer, while explicitly notifying the user of the missing visual input.
Figure A1: Case analysis of various MLLMs using ReasonMap (Case N1). For reasoning models, the reasoning process is explicitly marked with <think> and </think> tags. We highlight error contents in the answers with red and correct contents in green.
Figure A2: Case analysis of various MLLMs using ReasonMap (Case N2). For reasoning models, the reasoning process is explicitly marked with <think> and </think> tags. We highlight error contents in the answers with red and correct contents in green.
Figure A3: Case analysis of various MLLMs using ReasonMap (Case N3). For reasoning models, the reasoning process is explicitly marked with <think> and </think> tags. We highlight error contents in the answers with red and correct contents in green.
Appendix D Further Discussions
D.1 High-Resolution Image Preprocessing.
We compare how different Multimodal Large Language Models (MLLMs) handle high-resolution image inputs in Table A2. Specifically, we examine three key components in their preprocessing pipelines: dynamic resolution handling, positional encoding, and token compression.
Table A2: Comparison of high-resolution image preprocessing strategies across different MLLMs. We use “--” to denote unspecified or unclear content.
Model Dynamic Resolution Handling Positional Encoding Token Compression
Qwen2.5-VL series [2] 2D-RoPE ($2\times 2$ Concat + MLP)
QVQ-72B-Preview [20] 2D-RoPE ($2\times 2$ Concat + MLP)
InternVL3 series [3] V2PE (Unshuffle + MLP)
Kimi-VL series [16] 2D-RoPE (Shuffle + MLP)
Skywork-R1V-38B [17, 18] -
Gemini [50] (Tiling+Resize) -
Doubao-1.5-Pro series [19] 2D-RoPE ($2\times 2$ Pooling + MLP)
  1. 1.
    Dynamic resolution handling refers to whether the model can directly accept images of arbitrary sizes without resizing or cropping. Most recent models support native resolution processing, enabling them to preserve fine-grained spatial information. In contrast, models like Gemini [50] rely on image tiling and resizing to fit fixed input constraints.
  2. 2.
    Positional encoding helps the model retain spatial structure among visual tokens. Common strategies include 2D Rotary Positional Encoding (2D-RoPE) [66], as seen in Qwen2.5-VL [2] and Doubao [19], or flexible alternatives like V2PE [67] in InternVL3 [3]. Some models (e.g., Gemini, Skywork-R1V [17, 18]) do not explicitly disclose their positional encoding scheme, which we mark as “–” in the table.
  3. 3.
    Token compression aims to reduce the number of visual tokens for more efficient processing. Different models adopt different strategies: Qwen2.5-VL and QVQ [20] compress tokens via $2\times 2$ patch concatenation followed by an MLP; InternVL3 [3] and Kimi-VL [16] utilize spatial transformations like pixel unshuffle or shuffle, also followed by MLPs; Doubao averages over $2\times 2$ patches before projection. Models without token compression may incur higher memory and computation costs when processing high-resolution inputs.
D.2 Limitations and Future Work
While ReasonMap provides a carefully curated benchmark for evaluating fine-grained visual reasoning with high-resolution transit maps, we acknowledge that it represents only one type of structured visual diagram. As such, caution should be taken when generalizing observations to other domains that involve different types of visual content or reasoning styles. Additionally, although efforts were made to ensure diversity across cities and languages, the current version may not fully capture all geographic or linguistic variations. Future iterations could further expand coverage and explore additional forms of reasoning to enhance generality.
D.3 Broader Impact
Advancing the capabilities of MLLMs in fine-grained visual reasoning has the potential to benefit a wide range of real-world applications, including navigation systems, urban planning tools, and assistive technologies for visually impaired individuals. By offering a structured and rigorous benchmark, ReasonMap encourages the development of MLLMs that can more effectively interpret complex visual artifacts and perform spatial reasoning. This could contribute to the long-term goal of building intelligent agents that interact more naturally and safely with human environments. Furthermore, the dataset’s emphasis on high-resolution, globally sourced transit maps promotes research that is inclusive of diverse visual formats and geographic contexts. We hope ReasonMap can serve as a step toward more transparent, robust, and generalizable multimodal systems.
Appendix E License and Consent Information
E.1 Ethical Considerations
All experiments are conducted on ReasonMap, which is built using publicly available transit maps collected in compliance with relevant licenses and usage terms. The maps are selected to ensure geographic diversity and legal validity. Upon code release, we provide the source of each map for further reference. ReasonMap is intended solely for academic research on fine-grained visual understanding and spatial reasoning in MLLMs. It does not redistribute any copyrighted map images. All annotations are based on public information, contain no personal data, and are created under academic oversight. The benchmark is not intended for safety-critical use. We take care to ensure fairness, legal compliance, and responsible data handling.
E.2 Public Implementation
We benchmark the visual understanding and reasoning performance on ReasonMap across a diverse set of publicly available MLLMs:
To ensure fair and reproducible evaluation, we implement all inference procedures by adhering closely to the official documentation and recommended practices of each model. The code is released under the MIT License to support transparency and reproducibility. Additionally, we provide detailed usage instructions on the project website to ensure easy access and reproducibility for future users.